# Analyzing Multifiltering Functions Using Multiparameter Discrete Morse Theory

### v1

@@ 6.2. Pareto primary simplices
We end this section by discussing the particular case where $f:K\rightarrow{\mathbb{R}}^{k}$ is the max-extension of a component-wise injective map defined on the set $K_{0}$ of vertices of $K$. Under these conditions, it was shown by Allili et al. [4] that there exists a subset of simplices $S\subseteq K$, called primary simplices, such the lower stars $\operatorname{Low}_{f}(\sigma)=\{\tau\geq\sigma\ |\ f(\tau)\preceq f(\sigma)\}$ of primary simplices $\sigma\in S$ partition $K$. Also, it was noted by Scaramuccia et al. [61] that these lower stars correspond to the level sets $L_{u}$ of $f$. Indeed, since $f$ is component-wise injective on $K_{0}$ and $f(\tau)=\max_{v\in\tau}f(v)$ for all $\tau\in K$, then for every $u=(u_{1},...,u_{k})\in f(K)$ and each $i=1,...,k$, there exists a unique vertex $v_{i}\in K_{0}$ such that $f_{i}(v_{i})=u_{i}$. We could prove that $v_{1},v_{2},...,v_{k}$ are exactly the vertices of $\sigma=\bigcap_{\tau\in L_{u}}\tau$ (thus $\dim\sigma\leq k-1$) and that $\sigma$ is the unique simplex in $K$ such that $\operatorname{Low}_{f}(\sigma)=L_{u}$.
@@ 7.1. Space complexity reduction
Thus, we compare the numbers of critical simplices obtained using GenerateMDM with those of algorithms in [3, 4, 20]. Note that the algorithm presented in [20] does not compute a gradient field, but instead reduces the number of vertices and edges of the input dataset for while preserving the multipersistent homology of the space. Also, since ComputeDiscreteGradient [61] produces an output equivalent to that of Matching, we do not compare GenerateMDM to ComputeDiscreteGradient.
For each algorithm, we consider four different triangular meshes, available online in the GNU Triangulated Surface (GTS) Library [58]. As an input admissible map, we use the max-extension of $f:K_{0}\rightarrow{\mathbb{R}}^{2}$ such that $f(v)=(|x|,|y|)$ for each vertex $v$ of coordinates $(x,y,z)$. Also, the index mapping described at the beginning of Section 4.1 is used in GenerateMDM.
Table 1. Pourcentage of critical simplices output by Algorithm 1 and algorithms in [3, 4, 20]
Dataset tie space_shuttle x_wing space_station
$|K_{0}|$ 2014 2376 3099 5749
% 8.6 4.1 4.9 29.4
%[4] 27.5 9.5 19.8 30.8
%[3] 11.3 5.1 5.6 32.7
%[20] 29.2 11.0 18.4 33.7
$|K_{1}|$ 5944 6330 9190 15949
% 7.6 1.6 3.5 15.1
%[4] 20.1 3.8 13.4 16.0
%[3] 56.2 58.4 39.2 70.0
%[20] 13.9 5.2 9.2 17.4
$|K_{2}|$ 3827 3952 6076 10237
% 4.1 0.1 2.5 7.3
%[4] 14.1 0.4 9.9 8.0
%[3] 78.7 90.5 56.2 91.0
$|K|$ 11785 12658 18365 31935
% 6.7 1.6 3.4 15.2
%[4] 19.4 3.8 13.3 16.1
%[3] 55.9 58.4 39.2 70.0
The results are presented in Table 1. In the first block of the table, the first line shows the total number of vertices and the following three lines the percentage of simplices which are critical in the output of each algorithm. Analogous information is given in the two following blocks for edges and triangles, while the last block gives the total number of simplices for each mesh and the associated proportion of output critical simplices.
It is clear from Table 1 that, in terms of critical simplices produced, GenerateMDM outperforms all its predecessors. This is partially explained by the fact that the algorithms are tested on datasets for which the considered vertex map $v\mapsto(|x|,|y|)$ is not component-wise injective. As explained in Section 4.3, this means other algorithms which require this hypothesis to be verified have to preprocess the input function by perturbing it slightly, which can potentially induce factitious critical simplices.
@@ 7.2. Dependence on the indexing map and triangulation
To do so, we consider the following particular constructions of $I$. Let $K$ be a simplicial complex embedded in ${\mathbb{R}}^{3}$ with vertices $K_{0}=\{v_{0},...,v_{n}\}$. Note $(x_{i},y_{i},z_{i})$ the coordinates of each $v_{i}\in K_{0}$. We can choose to label the vertices so that $z_{i}<z_{j}\Rightarrow i<j$, meaning that the labels of the vertices increase as we move along the $x$-axis. Then, each $\sigma\in K$ may be represented by the labels of its vertices in decreasing order, and we can define $I:K\rightarrow{\mathbb{N}}$ as the index map obtained by ordering lexicographically $K$. We could show that $I$ is an admissible indexing map such that $\max_{v_{i}\in\tau}z_{i}<\max_{v_{j}\in\sigma}z_{j}\Rightarrow I(\tau)<I(\sigma)$, so the value of $I$ globally increases in the direction of the $z$-axis. We call $z$-increasing such an admissible index map. Alternatively, we can choose to label the vertices so that $z_{i}<z_{j}\Rightarrow i>j$ in order to obtain a $z$-decreasing indexing map such that $\max_{v_{i}\in\tau}z_{i}<\max_{v_{j}\in\sigma}z_{j}\Rightarrow I(\tau)>I(\sigma)$. Similarly, we can define indexing maps which are increasing or decreasing along the $x$ or $y$ axis.
@@
Figure 8. In 8a, the Pareto set $\mathcal{P}_{f}$, where $f$ is the max-extension of the vertex map $v\mapsto(x,y)$. The primary Pareto simplices are in blue, those nonprimary are in red. In 8b, the output of GenerateMDM when given $f$ and a $z$-increasing index map as input. In 8c, the output of GenerateMDM when given $f$ and a $z$-decreasing index map as input.
@@
Figure 9. Two outputs of GenerateMDM when considering a triangulated sphere $K$ and $f=0$ as input along with two different indexing maps $I$. We have in 9a the output for a $x$-increasing map $I$ and in 9b for a $y$-increasing $I$.
@@
Figure 10. Critical simplices of four different outputs of GenerateMDM. In each subfigure, the max-extension of the vertex map $v\mapsto(x,y)$ is considered as input function.
a
b
c
d
Figure 11. The Pareto set of the max-extension of the vertex map $v\mapsto(x,y)$ for four different triangulations of a sphere. The primary Pareto simplices are in blue, those nonprimary are in red. The subfigures correspond to those of Figure 10.
Furthermore, the way a given space is triangulated may influence significantly the critical simplices produced. To verify this, we use multiple meshes of the sphere generated with the Stripy library [55]. As shown in Figure 10, different triangulations yield visually dissimilar critical sets. That being said, we can see, in Figure 11, that this is due to the shape of the Pareto set of the input function on the given triangulation, and not to the algorithm itself.
@@ 7.3. Critical components
We now see how the definition of critical components proposed in Section 3.3 performs in practice. Recall that, for a mdm function $g$, we partition its critical simplices using the equivalence relation $\sim_{g}$ defined as the transitive closure of $R_{g}$, such that $\sigma R_{g}\tau$ when
  1. $(1)$
@@ 2. $(2)$
Condition $(1)$ ensures that $\sigma$ and $\tau$ can enter the multifiltration induced by $g$ at a same step, while condition $(2)$ implies that $\sigma$ and $\tau$ are connected and can interact with each other homologically.
@@
The previous examples and Figure 7 make it clear that, if we consider $K$ to be a triangulated sphere and $g:K\rightarrow{\mathbb{R}}^{2}$ to be a mdm function approximating the max-extension of the vertex map $v\mapsto(x,y)$, we can expect the critical simplices of $g$ to form two distinct critical components along the equator of $K$. As illustrated in Figure 12, $\sim_{g}$ yields the desired result for the given triangulation. Namely, we could characterize the green component in Figure 12a as a Pareto minimal component and the red one in Figure 12b as Pareto maximal.
@@
Figure 13. Critical components of $g$, where $g$ is the mdm function output by GenerateMDM when given as input the max-extension of the vertex map $v\mapsto(x,y)$. In 13a, the critical components are those with respect to $\sim_{g}$. In 13b, we consider $\sim_{g}^{\prime}$.
However, for some triangulations, the critical components induced by $\sim_{g}$ are not as desired. In Figure 13a, for the given triangulation (which is the same as that in Figures 10c and 11c), we see that some simplices which should belong in a same critical component are separated in a few different components.
@@
Figure 14. In 14a, a mdm function $g$ and its gradient field output by GenerateMDM, as shown in Figure 4. The critical components of $g$ with respect to $\sim_{g}$ are represented in green, pink, blue and orange. In 14b, the admissible map $f$ used to generate $g$. The critical components of $g$ with respect to $\sim_{g}^{\prime}$ are represented in green, pink and orange.
@@
To resolve this issue, when $g$ is a mdm function output by GenerateMDM with a multifiltering function $f$ as input, we can substitute condition $(1)$ of the definition of $R_{g}$ given previously by
@@
Figure 16. In 16a, critical components of $g$ with respect to $\sim_{g}^{\prime}$, where $g$ is the mdm function output by GenerateMDM when given the max-extension $f$ of the vertex map $v\mapsto(x,y)$ as input. There are three components, represented in blue, purple and red. The arrows show a gradient path connecting the critical simplices of the blue component along the $x=0$ plane. In 16b, a view of the torus zoomed in on the blue critical component.
@@
Note that this could not happen if the input map $f$ was the max-extension of a component-wise injective vertex map. Nevertheless, to ensure the concept of critical components yields desirable results in our more general setting, we also substitute condition $(2)$ of the definition of $R_{g}$ in order to obtain the following definition.
@@
The relation $R_{f}$ between critical simplices follows the idea of $R_{g}^{\prime}$. The first condition ensures that $\sigma$ and $\tau$ can enter the multifiltration induced by $f$ at a same step and the second implies that $\sigma$ and $\tau$ are connected in some way. The difference is in the idea of connectedness: the relation $R_{g}^{\prime}$ uses the gradient field of $g$ to determine if $\sigma$ and $\tau$ are connected, while $R_{f}$ uses their level sets with respect to $f$. Furthermore, when $g$ is perfect relatively to $f$, we can deduce from condition $(2)$ of Definition 7.1 that, if two critical simplices $\sigma$ and $\tau$ are such that $\sigma\sim_{f}\tau$, then they are connected by a path entirely contained in $\mathcal{P}_{f}$. Hence, for a relative-perfect mdm function $g$, each of its critical components with respect to $\sim_{f}$ is a subset of some connected component of $\mathcal{P}_{f}$.
@@
Figure 17. In 17a, the different colors show the connected components of the Pareto set $\mathcal{P}_{f}$, where $f$ is the max-extension of $v\mapsto(y,z)$ on the seashell dataset. In 17b, the corresponding critical components of $g$ with respect to $\sim_{f}$, where $g$ is the mdm function output by GenerateMDM when given $f$ as input.
In some cases, the critical components of the mdm function $g$ with respect to $\sim_{f}$ are in correspondence with the connected components of $\mathcal{P}_{f}$. For example, we can verify that if $g$ is defined on a sphere as in Figure 13, its critical components with respect to $\sim_{f}$ are the same as those with respect to $\sim_{g}^{\prime}$. Also, if it is defined on a torus as in Figures 15 and 16, it has four components with respect to $\sim_{f}$, as desired. Moreover, if $f$ is the max-extension of $v\mapsto(y,z)$ on the seashell triangulation obtained from the GTS library [58], as in Figure 17, we can verify that all critical components of the generated mdm function $g$ with respect to $\sim_{f}$ are in correspondence with the connected components of $\mathcal{P}_{f}$.
@@
Figure 18. In 18a, the different colors show the connected components of the Pareto set $\mathcal{P}_{f}$, where $f$ is the max-extension of $v\mapsto(x,z)$ on the head dataset from the GTS library. In 18b, the critical components of $g$ with respect to $\sim_{f}$, where $g$ is the mdm function output by GenerateMDM when given $f$ as input.
Nevertheless, in many cases, some of the connected components of $\mathcal{P}_{f}$ contain multiple critical components of $g$ with respect to $\sim_{f}$. See, for instance, functions $f$ and $g$ as defined in Figure 18. When this happens, we can interpret the critical components of $g$ as follows. Consider two critical simplices $\sigma$ and $\tau$ of $g$ and let $C_{\sigma}$ and $C_{\tau}$ be their connected components in $L_{f(\sigma)}$ and $L_{f(\tau)}$ respectively. Also, suppose $\sigma$ and $\tau$ are such that there exists $\sigma^{\prime}\in C_{\sigma}$ and $\tau^{\prime}\in C_{\tau}$ such that $\sigma^{\prime}\leq\tau^{\prime}$, meaning that condition $(2)$ of Definition 7.1 is satisfied. Moreover, assuming $g$ is perfect relatively to $f$, we have that $\sigma,\tau\in\mathcal{P}_{f}$ and belong in the same connected component. If $\sigma\not\sim_{f}\tau$, necessarily, condition $(1)$ of Definition 7.1 is not verified, so $f_{i}(\sigma)\neq f_{i}(\tau)$ for each $i=1,...,k$ and, because $\sigma^{\prime}\leq\tau^{\prime}$ implies $f(\sigma)=f(\sigma^{\prime})\preceq f(\tau^{\prime})=f(\tau)$, we deduce that $f_{i}(\sigma)<f_{i}(\tau)$ for each $i$. Thus, although $\sigma$ and $\tau$ belong in the same connected component of $\mathcal{P}_{f}$, we see that $\sigma$ always appear before $\tau$ in the multifiltration induced by $f$, so they cannot possibly enter it at a same step. In other words, $\sigma$ and $\tau$ correspond to two distinct features of the multifiltration, which explains why they belong to two different critical components with respect to $\sim_{f}$.

### v3

@@ 5.2. Pareto primary simplices
We end this section by discussing the particular case where $f:K\rightarrow{\mathbb{R}}^{k}$ is the max-extension of a component-wise injective map defined on the set $K_{0}$ of vertices of $K$. Under these conditions, it was shown in [4] that there exists a subset of simplices $S\subseteq K$, called primary simplices, such the lower stars $\operatorname{Low}_{f}(\sigma)=\{\tau\geq\sigma\ |\ f(\tau)\preceq f(\sigma)\}$ of primary simplices $\sigma\in S$ partition $K$. Also, it was noted in [63] that these lower stars correspond to the level sets $L_{u}$ of $f$. Indeed, since $f$ is component-wise injective on $K_{0}$ and $f_{i}(\tau)=\max_{v\in\tau}f_{i}(v)$ for all $\tau\in K$, then for every $u=(u_{1},...,u_{k})\in f(K)$ and each $i=1,...,k$, there exists a unique vertex $v_{i}\in K_{0}$ such that $f_{i}(v_{i})=u_{i}$. We could prove that $v_{1},v_{2},...,v_{k}$ are exactly the vertices of $\sigma=\bigcap_{\tau\in L_{u}}\tau$ (thus $\dim\sigma\leq k-1$) and that $\sigma$ is the unique simplex in $K$ such that $\operatorname{Low}_{f}(\sigma)=L_{u}$.
@@ 6.1. Space complexity reduction
Thus, we compare the numbers of critical simplices output by GenerateMDM with those of algorithms in [3, 4, 20]. Note that the algorithm presented in [20] does not compute a gradient field, but instead reduces the number of vertices and edges of the input dataset for while preserving the multipersistent homology of the space. Also, since ComputeDiscreteGradient [63] produces an output equivalent to that of Matching, we do not compare GenerateMDM to ComputeDiscreteGradient.
For each algorithm, we consider four different triangular meshes, available online in the GNU Triangulated Surface (GTS) Library [60]. As input admissible map, we use the max-extension of $f:K_{0}\rightarrow{\mathbb{R}}^{2}$ such that $f(v)=(|x|,|y|)$ for each vertex $v$ of coordinates $(x,y,z)$. Also, the index mapping described at the beginning of Section 4.1 is used in GenerateMDM.
Table 1. Proportion of critical simplices output by Algorithm 1 and algorithms in [3, 4, 20] given the vertex map $v\mapsto(|x|,|y|)$ and different datasets
Dataset Simplices Critical simplices (%)
Type Number Alg. 1 [4] [3] [20]
tie Vertices 2014 8.6 27.5 11.3 29.2
Edges 5944 7.6 20.1 56.2 13.9
Triangles 3827 4.1 14.1 78.7 -
Total 11785 6.7 19.4 55.9 -
space_shuttle Vertices 2376 4.1 9.5 5.1 11.0
Edges 6330 1.6 3.8 58.4 5.2
Triangles 3952 0.1 0.4 90.5 -
Total 12658 1.6 3.8 58.4 -
x_wing Vertices 3099 4.9 19.8 5.6 18.4
Edges 9190 3.5 13.4 39.2 9.2
Triangles 6076 2.5 9.9 56.2 -
Total 18365 3.4 13.3 39.2 -
space_station Vertices 5749 29.4 30.8 32.7 33.7
Edges 15949 15.1 16.0 70.0 17.4
Triangles 10237 7.3 8.0 91.0 -
Total 31935 15.2 16.1 70.0 -
The results are presented in Table 1 where, for each dataset, the number of simplices of each dimension is shown and the proportion of critical simplices output by each algorithm is presented by dimension. It is clear that, in terms of critical simplices produced, algorithm GenerateMDM outperforms all its predecessors. This is partially explained by the fact that the algorithms are tested on datasets for which the considered vertex map $v\mapsto(|x|,|y|)$ is not component-wise injective. As explained in Section 4.4, this means other algorithms which require this hypothesis to be verified have to preprocess the input function by perturbing it slightly, which can potentially induce factitious critical simplices.
@@ 6.2. Dependence on the indexing map and triangulation
To do so, we consider the following particular constructions of $I$. Let $K$ be a simplicial complex embedded in ${\mathbb{R}}^{3}$ with vertices $K_{0}=\{v_{0},...,v_{n}\}$. Note $(x_{i},y_{i},z_{i})$ the coordinates of each $v_{i}\in K_{0}$. We can choose to label the vertices so that $z_{i}<z_{j}\Rightarrow i<j$, meaning that the labels of the vertices increase as we move along the $z$-axis. Then, each $\sigma\in K$ may be represented by the labels of its vertices in decreasing order, and we can define $I:K\rightarrow{\mathbb{N}}$ as the index map obtained by ordering lexicographically $K$. We could show that $I$ is an admissible indexing map such that $\max_{v_{i}\in\tau}z_{i}<\max_{v_{j}\in\sigma}z_{j}\Rightarrow I(\tau)<I(\sigma)$, so the value of $I$ globally increases in the direction of the $z$-axis. We call $z$-increasing such an admissible index map. Alternatively, we can choose to label the vertices so that $z_{i}<z_{j}\Rightarrow i>j$ in order to obtain a $z$-decreasing indexing map such that $\max_{v_{i}\in\tau}z_{i}<\max_{v_{j}\in\sigma}z_{j}\Rightarrow I(\tau)>I(\sigma)$. Similarly, we can define indexing maps which are increasing or decreasing along the $x$ or $y$ axis.
@@
Figure 7. In 7a, the Pareto set $\mathcal{P}_{f}$, where $f$ is the max-extension of the vertex map $v\mapsto(x,y)$. The primary Pareto simplices are in blue, those nonprimary are in red. In 7b, the output of GenerateMDM when given $f$ and a $z$-increasing index map as input. In 7c, the output of GenerateMDM when given $f$ and a $z$-decreasing index map as input.
@@
Figure 8. Two outputs of GenerateMDM when considering a triangulated sphere $K$ and $f=0$ as input along with two different indexing maps $I$. We have in 8a the output for a $x$-increasing map $I$ and in 8b for a $y$-increasing $I$.
@@
Figure 9. In 9a-9b, critical simplices for different outputs of GenerateMDM. In each subfigure, the max-extension of the vertex map $v\mapsto(x,y)$ is considered as input function, but the chosen triangulations of the sphere differ. In 9c-9d, the corresponding Pareto sets. The primary Pareto simplices are in blue, those nonprimary are in red.
Furthermore, the way a given space is triangulated may influence significantly the critical simplices produced. To verify this, we use multiple meshes of the sphere generated with the Stripy library [57]. As shown in Figure 9, different triangulations yield visually dissimilar critical sets. That being said, we can see that this is due to the shape of the Pareto set of the input function on the given triangulation, and not to the algorithm itself.
@@ 6.3. Critical components
We now see how Definition 3.5 of critical components, first introduced in [14], performs in practice. Recall that, for a mdm function $g$, we partition its critical simplices using the equivalence relation $\sim_{g}$ defined as the transitive closure of $R_{g}$, such that $\sigma R_{g}\tau$ when
  1. $\mathrm{(1)}$
@@ 2. $\mathrm{(2)}$
Condition $\mathrm{(1)}$ ensures that $\sigma$ and $\tau$ can enter the multifiltration induced by $g$ at a same step, while condition $\mathrm{(2)}$ implies that $\sigma$ and $\tau$ are connected and can interact with each other homologically.
@@
The previous examples and Figure 6 make it clear that, if we consider $K$ to be a triangulated sphere and $g:K\rightarrow{\mathbb{R}}^{2}$ to be a mdm function approximating the max-extension of the vertex map $v\mapsto(x,y)$, we can expect the critical simplices of $g$ to form two distinct critical components along the equator of $K$. As illustrated in Figure 10, $\sim_{g}$ yields the desired result for the given triangulation. Namely, we could characterize the blue component in Figure 10a as a Pareto minimal component and the red one in Figure 10b as Pareto maximal.
@@
Figure 11. Critical components of $g$, where $g$ is the mdm function output by GenerateMDM when given as input the max-extension of the vertex map $v\mapsto(x,y)$. In 11a, the critical components are those with respect to $\sim_{g}$. In 11b, we consider $\sim_{g}^{\prime}$.
However, for some triangulations, the critical components induced by $\sim_{g}$ are not as desired. In Figure 11a, for the given triangulation (which is the same as that in Figures 9b and 9d), we see that some simplices which should belong in a same critical component are separated in a few different components.
@@
Figure 12. In 12a, a mdm function $g$ and its gradient field output by GenerateMDM, as shown in Figure 4. The critical components of $g$ with respect to $\sim_{g}$ are represented in green, pink, blue and orange. In 12b, the admissible map $f$ used to generate $g$. The critical components of $g$ with respect to $\sim_{g}^{\prime}$ are represented in green, pink and orange.
@@
To resolve this issue, when $g$ is a mdm function output by GenerateMDM with a multifiltering function $f$ as input, we can substitute condition $\mathrm{(1)}$ of the definition of $R_{g}$ given previously by
@@
Figure 14. In 14a, critical components of $g$ with respect to $\sim_{g}^{\prime}$, where $g$ is the mdm function output by GenerateMDM when given the max-extension $f$ of the vertex map $v\mapsto(x,y)$ as input. There are three components, represented in blue, purple and red. The arrows show a gradient path connecting the critical simplices of the blue component along the $x=0$ plane. In 14b, a view of the torus zoomed in on the blue critical component.
@@
Note that this could not happen if the input map $f$ was the max-extension of a component-wise injective vertex map. Nevertheless, to ensure the concept of critical components yields desirable results in our more general setting, we also substitute condition $\mathrm{(2)}$ of the definition of $R_{g}$ in order to obtain the following definition.
@@
The relation $R_{f}$ between critical simplices follows the idea of $R_{g}^{\prime}$. The first condition ensures that $\sigma$ and $\tau$ can enter the multifiltration induced by $f$ at a same step and the second implies that $\sigma$ and $\tau$ are connected in some way. The difference is in the idea of connectedness: the relation $R_{g}^{\prime}$ uses the gradient field of $g$ to determine if $\sigma$ and $\tau$ are connected, while $R_{f}$ uses their level sets with respect to $f$. Furthermore, when $g$ is perfect relatively to $f$, we can deduce from condition $\mathrm{(2)}$ of Definition 6.1 that, if two critical simplices $\sigma$ and $\tau$ are such that $\sigma\sim_{f}\tau$, then they are connected by a path entirely contained in $\mathcal{P}_{f}$. Hence, for a relative-perfect mdm function $g$, each of its critical components with respect to $\sim_{f}$ is a subset of some connected component of $\mathcal{P}_{f}$.
@@
Figure 15. In 15a, the different colors show the connected components of the Pareto set $\mathcal{P}_{f}$, where $f$ is the max-extension of $v\mapsto(y,z)$ on the seashell dataset. In 15b, the corresponding critical components of $g$ with respect to $\sim_{f}$, where $g$ is the mdm function output by GenerateMDM when given $f$ as input.
In some cases, the critical components of the mdm function $g$ with respect to $\sim_{f}$ are in correspondence with the connected components of $\mathcal{P}_{f}$. For example, we can verify that if $g$ is defined on a sphere as in Figure 11, its critical components with respect to $\sim_{f}$ are the same as those with respect to $\sim_{g}^{\prime}$. Also, if it is defined on a torus as in Figures 13 and 14, it has four components with respect to $\sim_{f}$, as desired. Moreover, if $f$ is the max-extension of $v\mapsto(y,z)$ on the seashell triangulation obtained from the GTS library [60], as in Figure 15, we can verify that all critical components of the generated mdm function $g$ with respect to $\sim_{f}$ are in correspondence with the connected components of $\mathcal{P}_{f}$.
@@
Figure 16. In 16a, the different colors show the connected components of the Pareto set $\mathcal{P}_{f}$, where $f$ is the max-extension of $v\mapsto(x,z)$ on the head dataset from the GTS library. In 16b, the critical components of $g$ with respect to $\sim_{f}$, where $g$ is the mdm function output by GenerateMDM when given $f$ as input.
Nevertheless, in many cases, some of the connected components of $\mathcal{P}_{f}$ contain multiple critical components of $g$ with respect to $\sim_{f}$. See, for instance, functions $f$ and $g$ as defined in Figure 16. When this happens, we can interpret the critical components of $g$ as follows. Consider two critical simplices $\sigma$ and $\tau$ of $g$ and let $C_{\sigma}$ and $C_{\tau}$ be their connected components in $L_{f(\sigma)}$ and $L_{f(\tau)}$ respectively. Also,we choose $\sigma$ and $\tau$ so that there exists $\sigma^{\prime}\in C_{\sigma}$ and $\tau^{\prime}\in C_{\tau}$ which satisfy $\sigma^{\prime}\leq\tau^{\prime}$, meaning that condition $\mathrm{(2)}$ of Definition 6.1 is verified. Moreover, assuming $g$ is perfect relatively to $f$, it follows from Corollary 5.4 that $\sigma,\tau\in\mathcal{P}_{f}$ and, since $\sigma^{\prime}\leq\tau^{\prime}$, we see that $\sigma$ and $\tau$ belong in the same connected component of $\mathcal{P}_{f}$. If $\sigma\not\sim_{f}\tau$, necessarily, condition $\mathrm{(1)}$ of Definition 6.1 is not satisfied, so $f_{i}(\sigma)\neq f_{i}(\tau)$ for each $i=1,...,k$ and, because $f(\sigma)=f(\sigma^{\prime})\preceq f(\tau^{\prime})=f(\tau)$, we deduce that $f_{i}(\sigma)<f_{i}(\tau)$ for each $i$. Thus, although $\sigma$ and $\tau$ belong in the same connected component of $\mathcal{P}_{f}$, we see that $\sigma$ always appear before $\tau$ in the multifiltration induced by $f$, so they cannot possibly enter it at a same step. In other words, $\sigma$ and $\tau$ correspond to two distinct features of the multifiltration, which explains why they belong to two different critical components with respect to $\sim_{f}$.